A lottery winner will receive $1 million at the end of each of the next ten years. What is the future value of her winnings at the time of her final payment (in millions of dollars), given that the interest rate is 9.5% per year? (Round your answer to three decimal places.)

$ million

--------------------------------------------------------------------------------------------------------------------------------------------------------------------------------

Your grandmother has been putting $1,000 into a savings account on every birthday since your first (that is, when you turned one). The account pays an interest rate of 2%. How much money (in $) will be in the account immediately after your grandmother makes the deposit on your twenty-first birthday? (Round your answer to the nearest cent.)

$

--------------------------------------------------------------------------------------------------------------------------------------------------------------------------------

You have a chance to buy an annuity that pays $450 at the beginning of each year for 3 years. You could earn 7.5% on your money in other investments with equal risk. What is the most (in $) you should pay for the annuity? (Round your answer to the nearest cent.)

$

--------------------------------------------------------------------------------------------------------------------------------------------------------------------------------

You want to buy a new sports car 3 years from now, and you plan to save $4,400 per year, beginning immediately. You will make 3 deposits in an account that pays 5.7% interest (the last deposit will be at the end of year 2). Under these assumptions, how much (in $) will you have 3 years from today? (Round your answer to the nearest cent.)

$
Answer #1
FVOrdinary Annuity = C*(((1 + i/100)^n -1)/(i/100))
C = Cash flow per period
i = interest rate
n = number of years
FV= 1*(((1+ 9.5/100)^10-1)/(9.5/100))
FV = 15.56
Using Calculator: press buttons "2ND"+"FV" then assign
PMT =-1
I/Y =9.5
N =10
PV = 0
CPT FV
Using Excel
=FV(rate,nper,pmt,pv,type)
=FV(9.5/(100),10,-1,,)
Please ask remaining parts seperately, questions are unrelated